# Hinge electronic structure of strained half-Heuslers

### v1

@@ I Introduction
The possibility of hinge states in gapless semimetals has also been recognized [21, 22, 23, 24] and the quest for specific materials where to study them is naturally important. Ref. [23] showed the existence of hinge states in models relevant for various candidate Dirac semimetals, such as Cd3As2 and KMgBi. Ref. [25] argued about the importance of hinge states in $\beta$-MoTe2 for the interpretation of its surface electronic structure while Ref. [26] reported evidence of hinge states in the related compound WTe2.
Recently, it has been suggessted the possibility of HOTI phases in the broad set of half-Heusler (hH) compounds [27]. Half-Heuslers are ternary compounds with the space group (SG) No. 216 [28, 29, 30]. They present at low energy a very similiar electronic structure to that of CdTe or HgTe, the latter one a reference example of nontrivial electronic topology since the first predictions [31, 32] and experimental confirmation [33, 34] of hosting a topological insulating phase under strain. The richness of HgTe goes hand in hand with the simplicity of its band structure, the hallmark of which is an inversion between the so-called $\Gamma_{6}$ and $\Gamma_{8}$ bands [35]. In the cubic phase, the $\Gamma_{8}$ bands exhibit a fourfould crossing at low energies which enforces a semimetallic phase. An appropriate reduction of the symmetry opens a gap, yielding a strong topological insulator (TI).
@@
The above-mentioned suggestion of HOTI phases in hHs is based on a study of a four-band model reasonably accurate to desribe the $\Gamma_{8}$ bands [27]. In this work, we investigate the problem by means of density-functional theory (DFT) calculations and of ab-initio derived tight-binding models. Our calculations show the existence of electronic states on certain hinges of hHs with either $\varepsilon_{\Gamma_{8}}>\varepsilon_{\Gamma_{6}}$ or with $\varepsilon_{\Gamma_{8}}<\varepsilon_{\Gamma_{6}}$. Importantly, we find these hinges states to be topologically trivial. Further, we show that they are critically affected by the topology of the bulk electronic structure due to trivial hybridization effects: topological surface states, when present, hybridize with the hinge modes which, as a consequence, avoid being exponentially localized around the hinges.
This paper is organized as follows. In Sec. II, we describe methodological aspects of this work. In Sec. III, we present the bulk and surface electronic structure of the lithium-based half-Heusler compounds LiBiZn and LiSbZn. In Sec. IV, we discuss relevant Hamiltonians that describe the low energy electronic properties of these compounds and in Sec. V we present hinge band structure calculations. Sec. VI contains our concluding remarks. The Appendix. A provides 8-band tight-binding Hamiltonians suitable to describe LiBiZn and LiSbZn.
@@ II Methods
Density functional theory (DFT) calculations were performed with the FPLO code [48] version 18.57, treating the spin-orbit coupling in the fully-relativistic four-component formalism and performing numerical Brillouin zone integrations with the tetrahedron method using a mesh of $30\times 30\times 30$ subdvisions. We relaxed the crystal structure by minimization of the total energy. For the simulation of uniaxial strain, we fixed the deformation of the lattice parameter $c$, and relaxed the perpendicular lattice parameters under the constraint $a=b$. To parametrize the deformation, we define $\delta=(c_{0}-c)/c_{0}$, where $c_{0}$ is the equilibrium lattice parameter.
Starting from the DFT results, tight-binding models were obtained by construction of Wannier functions with the projective method as implemented in FPLO [49]. Based on the analysis of the orbital-projected density of states, we consider an eight band model that includes Bi-6p (or Sb-5p) and Zn-4s orbitals. These tight-binding models were used for the study of the electronic structure at surfaces and hinges. For the former, we considered semi-infinite slabs using the the PYFPLO module of the FPLO code, while for the latter we solved finite systems using own codes based on Kwant [50] (more details are presented in Section V).
@@ III Lithium half-Heuslers
The crystal structure of a hH compound of formula $XYZ$ can be regarded as two displaced zinc-blende lattices formed by $XZ$ and $YZ$ pairs of atoms [Fig. 1(a)]. With this notation, $X$ and $Y$ occupy the cristallographically equivalent sites $4a$ and $4b$, forming a rock salt structure, while $Z$ is positioned at the site $4c$. We will consider the cases $X$$=$Li, $Y$$=$Bi or Sb and $Z$$=$Zn.
We choose Li-based hHs because, as we will show, their simplicity facilitate the construction of low-energy tight-binding models that accurately reproduce the density-functional calculations with a minimum effort in terms of number of bands. This allow us to tackle the problem of studying the hinge electronic structure using realistic tight-binding Hamiltonians.
While both LiSbZn and LiBiZn have originally been studied in the hexagonal phase [51, 52], the former has also been grown in the cubic phase [53] and the latter has recently been predicted to present a cubic phase as ground state [54]. We notice that within the same cubic space group, there are two variants defined by the crystallographic site occupied by the Zn atoms. In particular, in the family LiZnT (T$=$N, P, As), the Zn atoms are found in the $Y$ position. On the other hand, Ref. [53] has shown that the compound LiSbZn is different, since Zn occupies the $Z$ position [see Fig. 1(a)]. We have compared the total energy of these two variants for LiBiZn and LiSbZn, obtaining that the structure with Zn in the $Z$ site has lower energy. We notice that for both structures the Bi and Zn atoms — which provide the dominant contribution near the Fermi energy, as will be shown below — form a zinc-blende lattice and therefore, both variants have similar low-energy band structures.
@@
We obtain for the fully relaxed compounds the lattice parameters $a=6.584\,$Å   for LiBiZn and $a=6.350\,$Å   for LiSbZn. Figure 1 (c, d) shows the band structures and the orbital-projected density of states $D(\varepsilon)$ of LiBiZn and of LiSbZn, respectively. In both cases, Bi-$6p$ or Sb-$5p$ states dominate the electronic structure in the energy windows shown and exhibit a substantial hybridization with Zn-$4s$ and Li-$2s$. The main difference between compounds is the relative position of the $\Gamma_{6}$ and $\Gamma_{8}$ bands, which accommodate in total four electrons per unit cell. In LiSbZn, the energies of these bands obey $\varepsilon_{\Gamma_{8}}<\varepsilon_{\Gamma_{6}}$, so that the $\Gamma_{8}$ bands are completely filled and the compound is an insulator. In LiBiZn, the opposite arrangement leads to half-filled $\Gamma_{8}$ bands, enforcing a semimetallic state. In this case, a clear $s$-$p$ inversion is noticeable near $\Gamma$. In Appendix A we present an eight-band tight-binding model suitable to describe both compounds. An analysis of the model parameters indicates that the main difference between compounds is the larger local spin-orbit coupling in the Bi-based compound. Thus, these compounds having $\varepsilon_{\Gamma_{8}}>\varepsilon_{\Gamma_{6}}$ (LiBiZn) or $\varepsilon_{\Gamma_{8}}<\varepsilon_{\Gamma_{6}}$ (LiSbZn) provide a framework for the study of the hinge electronic structure of both band-inverted and normal systems having very similar chemical and physical properties (e.g. lattice parameters and orbital composition of low-energy bands).
Figure 2: (a) Band structure of LiBiZn under uniaxial strain $\delta=-0.02$. Thinner black curves are the bands without strain. (b) Band structure for $\delta=0.02$. The path chosen include one of the eight Weyl nodes found in the $k_{z}=0$ plane, of coordinates $\frac{2\pi}{a}(0.017,0.006,0)$.
@@
Figure 3: (a) [1,-1,0] Zn-terminated surface spectral density structure of LiBiZn under uniaxial strain $\delta=-0.02$. The panel on the right is a zoom of the white square in the left panel. (b) Analogous results for strain $\delta=0.02$. (c),(d) Zoom of the data in (a),(b), respectively. The path width in the zooms is 0.04×$\times 2\pi/a$.
To conclude this Section, we briefly present how these properties influence the surface spectral features. We focus on LiBiZn and consider the Zn-terminated [1,-1,0] surface, which is perpendicular to the direction in which strain is applied. Figures 3 (a) and (b) show results for $\delta=-0.02$ and $\delta=0.02$, respectively. The path chosen includes segments perpendicular ($K\Gamma$) or parallel ($\Gamma Z$) to the direction of the strain. Surface states with Dirac cones located at the Brillouin zone border ($K$ and $Z$) are observed for both signs of strain. We notice that for a LiBi-terminated surface, while the connectivity provided by the surface states between the projection of the bulk valence and conducting bands is the same, the Dirac cone at $Z$ or $K$ is pushed into the bulk projection (not shown). As shown in Figures 3 (c) and (d), the main role of strain is observed near $\Gamma$: the spectrum becomes gapped at $\Gamma$ and gapless at the projection of the Dirac cone (along $\Gamma Z$ and $\delta<0$) or of the Weyl nodes (along $K\Gamma$ and $\delta>0$).
IV Low energy models
@@
$\begin{split}\mathcal{H}(\mathbf{k})&=t\sum\limits_{j=1}^{2}\sin(k_{j}a)\gamma_{j}\\ &+\Big{[}-t_{3}\cos(k_{3}a)+m_{z}+t_{0}\sum\limits_{j=1}^{2}[1-\cos(k_{j}a)]\Big{]}\gamma_{3}\\ &\Delta_{1}[\cos(k_{2}a)-\cos(k_{1}a)]\gamma_{4}+\Delta_{2}\sin(k_{1}a)\sin(k_{2}a)\gamma_{5},\end{split}$ (1)
where $a$ is the lattice constant, $\mathbf{k}=(k_{1},k_{2},k_{3})$ is the momentum vector, and $\gamma_{j}$, $1\leq j\leq 5$, are the mutually anticommuting $4\times 4$ Hermitian matrices satisfying the algebra $\{{\gamma_{i},{\gamma_{j}}\}=2\delta_{ij}}$. Theses can be represented as $\gamma_{1}=\sigma_{3}\tau_{1}$, $\gamma_{2}=\sigma_{0}\tau_{2}$, $\gamma_{3}=\sigma_{0}\tau_{3}$, $\gamma_{4}=\sigma_{1}\tau_{1}$, $\gamma_{5}=\sigma_{2}\tau_{1}$, where $\sigma$ and $\tau$ Pauli matrices encode the spin and orbital degrees of freedom, respectively. As shown by Ref. [27], expansion of this model around $\Gamma$ can be mapped to the Luttinger Hamiltonian. The model describes well the low energy dispersion of band-inverted hHs, with the caveat that additional crystal symmetries make the crossing not isolated in actual hHs. Namely, within the model, the bands split in four at any momenta infinitesimally away from $\Gamma$ while in the material there are lines (e.g. $\Gamma$-$X$) where the bands split only in pairs [see Fig. 1 (c, d)]  [^1] See [47] for a demonstration that only materials in the SGs 207-214 can host fourfold crossings whose low energy description is exactly that of spin 3/2 quasiparticles..
The term $\delta=(-t_{3}\cos k_{3}a+m_{z})\gamma_{3}$ mimics the effects of strain along the $[001]$ direction  [^2] For consistenty with our DFT results, we define $\delta$ with an opposite sign to Ref. [27]. . Depending on the sign of $\delta$ at $k_{3}=0$, this term leads the model Eq. (1) to either a Dirac semimetal phase ($t_{3}>m_{z}$), or to a stack of quantum spin Hall insulating(3D) phases ($t_{3}<m_{z}$). Interestingly, these phases support higher-order topological modes along certain hinges [27].
@@
Figure 4: (a) Wilson loop in the $k_{z}=0$ plane for the four band model Eq. 1 with $\delta<0$. (b) Same for $\delta>0$. (c) Wannier center evolution in the plane defined by $\Gamma$ and two adjacent L points for LiZnBi under compressive strain. (d) Same as (c) in the case of tensile strain.
To illustrate the difference in topology, Figs. 4(a) and (b) show the Wilson loop computed in the $k_{3}=0$ plane for positive or negative strain, respectively, for the model Eq. (1). The different winding indicates a change in the value of the $\mathbb{Z}_{2}$ invariant for different signs of strain. On the other hand, the $\mathbb{Z}_{2}$ invariant in LiBiZn remains nontrivial independently of the sign of strain, as indicated by the evolution of the Wannier centers shown in Figs. 4(c) and (d) for compressive and tensile strain, respectively. Therefore, the electronic structure of both surfaces and hinges can be very different between a system described by the model Eq. (1) and hHs.
@@ V Hinge electronic structure
In this section, we present hinge band structure calculations for LiBiZn and LiSbZn. These are based on tight-binding Hamiltonians associated with Wannier functions constructed from DFT calculations. As explained above, a six-band model would be the minimal option to correctly capture the topology of band-inverted hHs. However, it is problematic to obtain such a model from a Wannierization processs consisting in the projection of the Khon-Sham wave functions onto pre-defined local orbitals, essentialy because the low-energy six bands have a strong mixture of Bi-6p and s-like orbitals (primarily from Zn but also in smaller amount from Li and Bi as well). This makes the definition of the local orbitals onto which project unknown. Therefore, at the expense of a numerically more demanding hinge bandstructure calculation, we shall work with 8-band models including all Bi 6p (or Sb 5p) and Zn 4s orbitals, as explained in Section II.
@@
Figures 5(d) and (e) present for both compounds the local density of states (LDOS) which could be measured, e.g., by scanning tunneling spectroscopy. In the case of LiBiZn, the LDOS is computed at $\varepsilon=0$, whereas $\varepsilon=0.2$ eV is chosen for LiSbZn. A relatively large LDOS at the Bi/Sb hinge can be observed in both compounds. In line with the results above, LiSbZn being both insulating in the bulk and at surface, the found hinge modes are exponentially localized at the hinge for all momenta and the LDOS accordingly decays very fast away from the hinge. On the other hand, a much larger spread of the hinge mode along the surface can be observed in LiBiZn. An analysis of the orbital composition indicates that hinge modes have a weight of $\sim 95$% on the Bi/Sb orbitals. Further, we find that uniaxial strain can affect the energy of the hinge modes. Changes in the order of 10 meV are found for a 2% lattice deformation.
Finally, we note that the energy and the real-space probability density of these boundary-localized modes depend on the crystal termination, i.e. the shape of the prism cross-section. We show in Appendix B an alternate choice of geometry, in which the cross-section is rectangular instead of rhombus-shaped. In this case, we observe that the modes become spread out over the entire Bi(Sb)-terminated surface.

### v2

# Electronic structure and hinge states of strained half-Heusler compounds LiSbZn and LiBiZn

@@ I Introduction
The possibility of hinge states in gapless semimetals has also been recognized [21, 22, 23, 24] and the quest for specific materials where to study them is naturally important. Reference [23] showed the existence of hinge states in models relevant for various candidate Dirac semimetals, such as Cd3As2 and KMgBi. Reference [25] argued about the importance of hinge states in $\beta$-MoTe2 for the interpretation of its surface electronic structure while Ref. [26] reported evidence of hinge states in the related compound WTe2.
Recently, it has been suggested the possibility of HOTI phases in the broad set of half-Heusler (hH) compounds [27]. Half-Heuslers are ternary compounds with the space group (SG) No. 216 [28, 29, 30]. They present at low energy a very similar electronic structure to that of CdTe or HgTe, the latter one a reference example of nontrivial electronic topology since the first predictions [31, 32] and experimental confirmation [33, 34] of hosting a topological insulating phase under strain. The richness of HgTe goes hand in hand with the simplicity of its band structure, the hallmark of which is an inversion between the so-called $\Gamma_{6}$ and $\Gamma_{8}$ bands [35]. In the cubic phase, the $\Gamma_{8}$ bands exhibit a fourfold crossing at low energies which enforces a semimetallic phase. An appropriate reduction of the symmetry opens a gap, yielding a strong topological insulator (TI).
@@
The above-mentioned suggestion of HOTI phases in hHs is based on a study of a four-band model reasonably accurate to describe the $\Gamma_{8}$ bands [27]. In this work, we investigate the problem by means of density-functional theory (DFT) calculations and of ab-initio derived tight-binding models. Our calculations show the existence of electronic states on certain hinges of hHs with either $\varepsilon_{\Gamma_{8}}>\varepsilon_{\Gamma_{6}}$ or with $\varepsilon_{\Gamma_{8}}<\varepsilon_{\Gamma_{6}}$. Importantly, we find these hinge states to be topologically trivial. Further, we show that they are critically affected by the topology of the bulk electronic structure due to trivial hybridization effects: Topological surface states, when present, hybridize with the hinge modes which, as a consequence, avoid being exponentially localized around the hinges.
This paper is organized as follows. In Sec. II, we describe methodological aspects of this work. In Sec. III, we present the bulk and surface electronic structure of the lithium-based half-Heusler compounds LiBiZn and LiSbZn. In Sec. IV, we discuss relevant Hamiltonians that describe the low energy electronic properties of these compounds and in Sec. V we present hinge band structure calculations. Sec. VI contains our concluding remarks. The Appendix. A provides 8-band tight-binding Hamiltonians suitable to describe LiBiZn and LiSbZn. Data and code that reproduce our main results are available on Zenodo [48].
@@ II Methods
Density functional theory (DFT) calculations were performed with the FPLO code [49] version 18.57, treating the spin-orbit coupling in the fully-relativistic four-component formalism and performing numerical Brillouin zone integrations with the tetrahedron method using a mesh of $30\times 30\times 30$ subdivisions. We relaxed the crystal structure by minimization of the total energy. For the simulation of uniaxial strain, we fixed the deformation of the lattice parameter $c$, and relaxed the perpendicular lattice parameters under the constraint $a=b$. To parametrize the deformation, we define $\delta=(c_{0}-c)/c_{0}$, where $c_{0}$ is the equilibrium lattice parameter. Our primary goal is to identify which phases are accessible under moderate uniaxial strain, specifically, whether a HOTI phase should be expected. For clarity in the band-structure analysis, we use representative strains of $\delta=\pm 2\%$. Although such values may be challenging to realize in this materials family, they sharpen the features central to our discussion. Importantly, our qualitative conclusions about the presence or absence of a HOTI phase do not depend on the precise magnitude of $\delta$.
Starting from the DFT results, tight-binding models were obtained by construction of Wannier functions with the projective method as implemented in FPLO [50]. Based on the analysis of the orbital-projected density of states, we consider an eight band model that includes Bi-6p (or Sb-5p) and Zn-4s orbitals. These tight-binding models were used for the study of the electronic structure at surfaces and hinges. For the former, we considered semi-infinite slabs using the PYFPLO module of the FPLO code, while for the latter we solved finite systems using own codes based on Kwant [51] (more details are presented in Section V).
@@ III Lithium half-Heuslers
The crystal structure of a hH compound of formula $XYZ$ can be regarded as two displaced zinc-blende lattices formed by $XZ$ and $YZ$ pairs of atoms [Fig. 1(a)]. With this notation, $X$ and $Y$ occupy the crystallographically equivalent sites $4a$ and $4b$, forming a rock salt structure, while $Z$ is positioned at the site $4c$. We will consider the cases XX==Li, YY==Bi or Sb and ZZ==Zn.
We choose Li-based hHs because, as we will show, their simplicity facilitate the construction of low-energy tight-binding models that accurately reproduce the density-functional calculations with a minimum effort in terms of number of bands. This allows us to tackle the problem of studying the hinge electronic structure using realistic tight-binding Hamiltonians.
While both LiSbZn and LiBiZn have originally been studied in the hexagonal phase [52, 53], the former has also been grown in the cubic phase [54] and the latter has recently been predicted to present a cubic phase as ground state [55]. We notice that within the same cubic space group, there are two variants defined by the crystallographic site occupied by the Zn atoms. In particular, in the family LiZnT (T==N, P, As), the Zn atoms are found in the $Y$ position. On the other hand, Ref. [54] has shown that the compound LiSbZn is different, since Zn occupies the $Z$ position [see Fig. 1(a)]. We have compared the total energy of these two variants for LiBiZn and LiSbZn, obtaining that the structure with Zn in the $Z$ site has lower energy. We notice that for both structures the Bi and Zn atoms — which provide the dominant contribution near the Fermi energy, as will be shown below — form a zinc-blende lattice and therefore, both variants have similar low-energy band structures.
@@
We obtain for the fully relaxed compounds the lattice parameters a=6.584a=6.584\,Å   for LiBiZn and a=6.350a=6.350\,Å   for LiSbZn. Figure 1 (c, d) shows the band structures and the orbital-projected density of states $D(\varepsilon)$ of LiBiZn and of LiSbZn, respectively. In both cases, Bi-$6p$ or Sb-$5p$ states dominate the electronic structure in the energy windows shown and exhibit a substantial hybridization with Zn-$4s$ and Li-$2s$. The main difference between compounds is the relative position of the $\Gamma_{6}$ and $\Gamma_{8}$ bands, which accommodate in total four electrons per unit cell. In LiSbZn, the energies of these bands obey $\varepsilon_{\Gamma_{8}}<\varepsilon_{\Gamma_{6}}$, so that the $\Gamma_{8}$ bands are completely filled and the compound is an insulator. In LiBiZn, the opposite arrangement leads to half-filled $\Gamma_{8}$ bands, enforcing a semimetallic state. In this case, a clear $s$-$p$ inversion is noticeable near $\Gamma$. In Appendix A we present an eight-band tight-binding model suitable to describe both compounds. An analysis of the model parameters indicates that the main difference between compounds is the larger local spin-orbit coupling in the Bi-based compound. Thus, these compounds having $\varepsilon_{\Gamma_{8}}>\varepsilon_{\Gamma_{6}}$ (LiBiZn) or $\varepsilon_{\Gamma_{8}}<\varepsilon_{\Gamma_{6}}$ (LiSbZn) provide a framework for the study of the hinge electronic structure of both band-inverted and normal systems having very similar chemical and physical properties (e.g. lattice parameters and orbital composition of low-energy bands).
Figure 2: (a) Band structure of LiBiZn under uniaxial strain $\delta=-0.02$. Thinner black curves are the bands without strain. (b) Band structure for $\delta=0.02$. The path chosen includes one of the eight Weyl nodes found in the $k_{z}=0$ plane, of coordinates $\frac{2\pi}{a}(0.017,0.006,0)$.
@@
[…]
IV Low-energy models
@@
$\begin{split}\mathcal{H}(\mathbf{k})&=t\sum\limits_{j=1}^{2}\sin(k_{j}a)\gamma_{j}\\ &+\Big[-t_{3}\cos(k_{3}a)+m_{z}+t_{0}\sum\limits_{j=1}^{2}[1-\cos(k_{j}a)]\Big]\gamma_{3}\\ &+\Delta_{1}[\cos(k_{2}a)-\cos(k_{1}a)]\gamma_{4}+\Delta_{2}\sin(k_{1}a)\sin(k_{2}a)\gamma_{5},\end{split}$ (1)
where $a$ is the lattice constant, $\mathbf{k}=(k_{1},k_{2},k_{3})$ is the momentum vector, and $\gamma_{j}$, $1\leq j\leq 5$, are the mutually anticommuting $4\times 4$ Hermitian matrices satisfying the algebra $\{{\gamma_{i},{\gamma_{j}}\}=2\delta_{ij}}$. These can be represented as $\gamma_{1}=\sigma_{3}\tau_{1}$, $\gamma_{2}=\sigma_{0}\tau_{2}$, $\gamma_{3}=\sigma_{0}\tau_{3}$, $\gamma_{4}=\sigma_{1}\tau_{1}$, $\gamma_{5}=\sigma_{2}\tau_{1}$, where $\sigma$ and $\tau$ Pauli matrices encode the spin and orbital degrees of freedom, respectively. As shown by Ref. [27], expansion of this model around $\Gamma$ can be mapped to the Luttinger Hamiltonian. The model describes well the low-energy dispersion of band-inverted hHs, up to differences stemming from distinct symmetries. In particular, for real coefficients, $\mathcal{H}(\mathbf{k})$ is invariant at each $\mathbf{k}$ under the antiunitary operator $\tilde{T}=\sigma_{2}\tau_{3}K$ (not the physical time-reversal), where $K$ denotes complex conjugation and $\tilde{T}^{2}=-1$. Consequently, at generic $\mathbf{k}$ the bands are doubly degenerate. In actual hHs, spatial inversion is broken and spin-orbit coupling lifts the spin degeneracy, with doubly degenerate bands persisting only along high-symmetry directions (e.g. $\Gamma$-$X$).
The term $\delta=(-t_{3}\cos k_{3}a+m_{z})\gamma_{3}$ mimics the effects of strain along the $[001]$ direction  [^1] For consistency with our DFT results, we define $\delta$ with an opposite sign to Ref. [27].. Depending on the sign of $\delta$ at $k_{3}=0$, this term leads the model Eq. (1) to either a Dirac semimetal phase ($t_{3}>m_{z}$), or to a stack of quantum spin Hall insulator (3D) phases ($t_{3}<m_{z}$). Interestingly, these phases support higher-order topological modes along certain hinges [27].
@@
Figure 4: (a) Wilson loop in the $k_{z}=0$ plane for the four band model Eq. (1) with $\delta<0$. (b) Same for $\delta>0$. (c) Wannier center evolution in the plane defined by $\Gamma$ and two adjacent L points for LiZnBi under compressive strain. (d) Same as (c) in the case of tensile strain.
To illustrate the difference in topology, Figs. 4(a) and (b) show the Wilson loop computed in the $k_{3}=0$ plane for positive or negative strain, respectively, for the model Eq. (1). The different winding indicates a change in the value of the $\mathbb{Z}_{2}$ invariant for different signs of strain. On the other hand, the $\mathbb{Z}_{2}$ invariant in LiBiZn remains nontrivial independently of the sign of strain, as indicated by the evolution of the Wannier centers shown in Figs. 4(c) and (d) for compressive and tensile strain, respectively. Therefore, the electronic structure of both surfaces and hinges can be very different between a system described by the model Eq. (1) and hHs. Particularly, in band-inverted hHs the first-order topology makes the surfaces gapless, so that these systems do not qualify as HOTIs.
@@ V Hinge electronic structure
In this section, we present hinge band structure calculations for LiBiZn and LiSbZn. These are based on tight-binding Hamiltonians associated with Wannier functions constructed from DFT calculations. As explained above, a six-band model would be the minimal option to correctly capture the topology of band-inverted hHs. However, it is problematic to obtain such a model from a Wannierization process consisting in the projection of the Kohn-Sham wave functions onto pre-defined local orbitals, essentially because the low-energy six bands have a strong mixture of Bi-6p and s-like orbitals (primarily from Zn but also in smaller amount from Li and Bi as well). This makes the definition of the local orbitals onto which project unknown. Therefore, at the expense of a numerically more demanding hinge bandstructure calculation, we shall work with 8-band models including all Bi 6p (or Sb 5p) and Zn 4s orbitals, as explained in Section II.
@@
Notice that the hinge modes in LiSbZn do not interpolate between states above and below the Fermi level. Thus we conclude that LiSbZn is not a HOTI. Furthermore, the energy dispersion illustrates that there is no symmetry requirement on the energy of these hinge states, such that they could in principle be shifted in energy away from the gap, producing a trivial bandstructure.
Figures 5(d) and (e) present for both compounds the local density of states (LDOS) which could be measured, e.g., by scanning tunneling spectroscopy. In the case of LiBiZn, the LDOS is computed at $\varepsilon=0$, whereas $\varepsilon=0.2$ eV is chosen for LiSbZn. A relatively large LDOS at the Bi/Sb hinge can be observed in both compounds. In line with the results above, LiSbZn being both insulating in the bulk and at surface, the found hinge modes are exponentially localized at the hinge for all momenta and the LDOS accordingly decays very fast away from the hinge. This can be quantified by tracking the weight function $w_{\text{H}}$ associated to the hinge states as a function of momentum. We find $w_{\text{H}}\simeq 0.36$ close to the center of the Brillouin zone, going up to values $w_{\text{H}}\simeq 0.71$ for larger $\vec{k}\cdot\vec{a}_{1}$ [see Fig. 5(c)]. Notice that the hinge states are present only at certain corners, further indicating that their physical origin is associated with the crystal structure termination at the hinge, rather than with a bulk topological property. On the other hand, a much larger spread of the hinge mode along the surface can be observed in LiBiZn, as indicated also by the momentum-resolved weight functions presented in Appendix B. An analysis of the orbital composition indicates that hinge modes have a weight of $\sim 95$% on the Bi/Sb orbitals. Further, we find that uniaxial strain can affect the energy of the hinge modes. Changes in the order of 10 meV are found for a 2% lattice deformation.
Finally, we note that the energy and the real-space probability density of these boundary-localized modes depend on the crystal termination, i.e. the shape of the prism cross-section. We show in Appendix C an alternate choice of geometry, in which the cross-section is rectangular instead of rhombus-shaped. In this case, we observe that the modes become spread out over the entire Bi(Sb)-terminated surface.